Hi Could you please explain the steps of how to solve this problem for me? It would be greatly appericated... The x is suppose to be over 5 like a fraction but I can't type it as a fraction.
-22=x - 7
5

Answer by checkley71(8403)

-22=X/5-7 now add 7 to each side of the equation we get -22+7=X/5-7+7 OR
-15=X/5 now multiplying each side by 5 we get 5*-15=5x/5 or -75=x
PROOF -22=-75/5-7 OR -22=-15-7 OR -22=-22